OpenStudy (anonymous):

A SECTOR OF A CIRCLE HAS AN AREA OF 29 SQAURE FEET. FIND THE CENTRAL ANGLE IN DEGREES THAT FORS THE SECTOR IF THE RADIUS IS 7.3 FEET

Vocaloid (vocaloid):

(x/360)*(area of circle) = (area of sector)

Vocaloid (vocaloid):

solve for x

OpenStudy (anonymous):

I did , but I got a decimal and I need the answer in radians

Vocaloid (vocaloid):

huh? the question asks for degrees

OpenStudy (anonymous):

yeah but I got a decimal

OpenStudy (anonymous):

im sorry, can you show me how to work it out

Vocaloid (vocaloid):

decimals are fine, you can just round

OpenStudy (anonymous):

I know but I need teh naswers in radians

OpenStudy (anonymous):

*the answers

Vocaloid (vocaloid):

why??? the question doesn't want radians, it wants degrees

OpenStudy (anonymous):

ok then, how do i get the answers in degrees

OpenStudy (anonymous):

i still have a decimal

OpenStudy (anonymous):

how to i get it into degrees?

OpenStudy (anonymous):

I ned it in degrees

OpenStudy (mathstudent55):

Are of a sector of a circle with central angle \(\theta\): \(\Large A_{sector} = \dfrac{\theta}{360^o}\pi r^2\) \(\Large A_{sector} = \dfrac{\theta}{2\pi ~rad}\pi r^2\)

OpenStudy (anonymous):

mathstudent55, will that get my answer in degrees

OpenStudy (mathstudent55):